The pressure exerted by an ideal gas is given
by
p = 1/3 mnv²
Using this expression deduce Avogadro's
law.
1
Expert's answer
2019-05-21T08:53:20-0400

according to the original formula

"p = n \frac{mv^{2}}{2}\frac{2}{3} = nkT"

respectivly

"N = \frac{pV}{kT}"


and this is Avogadro's law - in equal volumes of different gases, taken at the same temperature and pressure, contains the same number of molecules.